### v2

@@ 5.2 Cross-lingual Analysis
We translate all questions and responses from CARE that are written in Chinese and Arabic into English to enable cross-lingual evaluation.
Figure 7(a) shows that LMs generally perform better on culture-specific questions when prompted in the native languages than in English, but more notably for Chinese than for Arabic. This advantage is also especially obvious for the literacy category, while the gap narrows in categories such as cultural commonsense. A possible explanation is that English corpora contain fewer discussions about Chinese or Arab cultures.
We also examine how language choice impacts search engine results. As shown in Figure 7(b), cultural information on entities, opinions, and literacy tends to be of higher quality when retrieved in native languages. Interestingly, this trend reverses for questions related to cultural commonsense and social norms, such as “In China, does the ticket time indicate when the feature movie starts, or are there trailers played before the main feature?”, where English searches yield higher-quality content. This aligns with the potential documentation gap mentioned in §5.1, where native speakers often assume everyday cultural knowledge is implicit, whereas non-native speakers are more likely to actively seek out and document it. A similar pattern is observed for Arabic. For a small number of Arabic questions (10% opinion, 14% social norms, 14% commonsense, and 7% literacy), the search engine also failed to return any results. One such example is “A Jordanian friend mentioned that it’s polite not to open a gift in front of the giver. Is this true?”.
@@ 6 Conclusion
We introduce CARE, a human-written, multilingual resource comprising 2,580 culture-specific questions about Chinese and Arab cultures, along with responses that include human preferences. Through extensive experiments, we investigate that a small amount of high-quality cultural alignment data can improve existing LMs via preference optimization, across different model families and scales. This cultural-specific alignment offers benefits that go beyond generic human preference and proves effective under various prompting setups. Our analyses also reveal gaps in cultural awareness among LMs, native-speaking humans, and search engines. For example, cultural commonsense is often left implicit in the native language but is more explicitly mentioned in foreign-language content. By releasing CARE, we hope to encourage the development of more inclusive and culturally adaptive LMs.
@@
In this work, as we investigate how to improve LMs by aligning with multilingual multicultural human preference data, we also identified a couple of areas for further exploration. First, while filtering the Arabic portion of the Aya dataset, we notice a large number (3,831) of questions centered on Islamic knowledge and religious history. We exclude them, as religion-specific understanding in LMs falls outside the scope of our study that focuses on culture-specific phenomena and factual knowledge, as defined in §3.1. Nonetheless, future research on religious QA or alignment could adopt a similar approach to CARE by collecting preference data from credible religious sources.
Second, LMs may occasionally respond to culture-specific questions in a stereotypical or biased manner, stemming from a lack of cultural understanding. For instance, when asked “In China, what does the term “laowai” mean when referring to foreigners?”, LMs often incorrectly interpret the term as disrespectful. In reality, “laowai” is typically a neutral descriptor in Chinese, used to denote foreigners without negative connotations. While CARE includes examples that clarify such misunderstandings through human-annotated responses, future studies in cultural red-teaming can further investigate and address these failure modes.
@@ Acknowledgments
The authors would like to thank Nour Allah El Senary, Jad Matthew Bardawil, Chao Jiang, and Yuming Pan for their help in data annotation; Yao Dou and Xiaofeng Wu for their helpful feedback. We would also like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This work was funded in part through a Sony Faculty Innovation Award, and by the NSF under grant number IIS-2144493, IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 Annotation Details
To construct the question-answer pairs in CARE, we recruited two native Chinese speakers and two native Arabic speakers who are undergraduate students. The annotators were paid at the rate of 16 USD/hour and told that they were helping create and annotate data for improving the cultural awareness of LMs. The detailed annotation guideline for dataset construction is provided in Figures 16 to 20, and the instruction for preference annotation is provided in Figures 14. We also show our rank-and-rate web interface for collecting human cultural preference pairs in Figure 15.
A.2 Statistics
Table 5 shows the detailed statistics for Chinese and Arab culture in CARE across the cultural categories, stratified by the data source. The questions specific to Chinese culture are written in Chinese, while the questions specific to Arab culture are written in Arabic. The culturally-relevant samples obtained from the instruction-tuning dataset mostly fall within the cultural entities, cultural opinion, or literacy categories, while samples obtained from cultural knowledge bases provide more social norms and cultural commonsense data.
Table 6 shows the detailed statistics for other foreign cultures in CARE for each cultural category. These samples were obtained when filtering the instruction tuning datasets and cultural knowledge bases (§3.1). We wrote these foreign samples in both Chinese and Arabic versions and used them as part of our training set and in our analysis on the impact of the source culture in alignment (§4.5). We also manually collected more samples for social norms and cultural commonsense from online websites to cross 100 samples in each category. We ensured these foreign samples do not include anything relevant to Chinese or Arab cultures.
Table 7 shows the detailed statistics for Chinese and Arabic samples in CARE across the geographic scope. Most samples fall into nationwide category, while sub-nationwide samples provide insights about more detailed cultural information, continent-wide and worldwide samples assess understanding between several cultures.
@@ Appendix B Implementation Details
DPO Alignment.
We perform preference optimization using DPO for 3 epochs until the loss converges. Training is done with a batch size of 128 on 4-8 NVIDIA A40 GPUs. We tune the learning rate in the range $\{1e^{-7},5e^{-7},9e^{-7}\}$ and set beta for DPO as 0.1. The training involves full fine-tuning with 5 warmup steps and employs a linear learning rate scheduler.
@@
For open-sourced LMs, we run inference on 1 NVIDIA A40 GPU with the vLLM library [^2] https://docs.vllm.ai Kwon et al. (2023). We perform decoding by setting the following parameters {temperature=0.7, top_p=1}. We limit the context length by setting {max_model_len=2048}. We also limit the number of generated tokens by the models by setting {max_tokens=1024}. For the closed-source GPT-4o LM, we run inference with Azure OpenAI API.
@@
You are a native [Chinese/Arab] person, familiar with
[Chinese/Arab] culture and traditions.
@@
Figure 8: Prompt template for role-play inference. The LM is told to take on the persona of a native [Chinese/Arab] person who is familiar with the culture’s traditions, and then asked the culture-specific {Question}.
@@
Chinese Arabic Model Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide [Uncaptioned image] Gemma2-27B 6.50 7.16 8.19 6.47 4.71 6.73 4.88 6.29 [Uncaptioned image] Llama3.3-70B 6.20 6.92 7.68 5.57 5.14 6.16 4.31 5.50 [Uncaptioned image] Qwen2.5-72B 7.70 8.49 8.60 6.78 7.86 7.62 5.19 8.58 [Uncaptioned image] Mistral-Large 7.42 8.11 8.73 6.70 5.29 7.24 5.31 8.08 [Uncaptioned image] GPT-4o 8.59 8.74 8.72 7.11 7.43 8.15 7.79 8.92
Table 8: Performance comparison w.r.t. geographic scope. Scores are computed on the entire CARE data, including both native culture and foreign culture.
@@ Appendix D Additional Results
D.1 Performance by Geographic Scope
Table 8 shows the results achieved by different LMs when stratified by geographic scope. We find that the models struggle the most with sub-nationwide questions. Most models achieve the best scores on continent-wide or worldwide questions.
D.2 Performance on other cultural tasks
We also evaluate the performance of LMs after cultural alignment with CARE on other cultural tasks using out-of-domain culture-related data. Specifically, we use the Include benchmark Romanou et al. (2025), as it offers multilingual versions for each culture and employs a multi-choice format, enabling us to assess the robustness of our alignment. From Table 9, we observe that alignment with culture-specific human preference can also benefit the performance on the Include benchmark, indicating the generalization capability of CARE.
[Uncaptioned image] Gemma2-9B [Uncaptioned image] Llama3.1-8B [Uncaptioned image] Qwen2.5-7B [Uncaptioned image] Mistral-7B Vanilla CARE Vanilla CARE Vanilla CARE Vanilla CARE Include (zh) 57.79 59.44 51.00 54.68 75.22 75.59 40.36 40.73 Include (ar) 57.42 54.71 50.54 51.08 56.15 55.97 35.32 36.95
Table 9: Accuracy (%) on the Include benchmark for Chinese and Arabic subsets, before and after cultural alignment with CARE. Improvements in the Include benchmark can also be observed after alignment with CARE.

### v4

@@ 5.2 Cross-lingual Analysis
We translate all questions and answers in CARE that are written in Chinese, Arabic, and Japanese into English to enable cross-lingual evaluation. Figure 6(a) shows that LMs generally perform better on culture-specific questions when prompted in the native languages than in English, but more notably for Chinese than Arabic and Japanese. This advantage is also especially obvious for the literacy category, while the gap narrows in categories such as cultural commonsense.
We also examine how language choice impacts search engine results. As shown in Figure 6(b), cultural information on entities, opinions, and literacy tends to be of higher quality when retrieved in native languages. Interestingly, this trend reverses for questions related to cultural commonsense and social norms, such as “In China, does the ticket time indicate when the feature movie starts, or are there trailers played before the main feature?”, where English searches yield higher-quality content. This aligns with the potential documentation gap mentioned in §5.1, where native speakers often assume everyday cultural knowledge is implicit, but non-native speakers are more likely to seek out information or document it.
@@ 6 Conclusion
We introduce CARE, a multilingual, human-written resource comprising 3,490 culture-specific questions about Chinese, Arab, and Japanese cultures, along with human-rated responses. Through extensive experiments, we investigate that a small amount of high-quality cultural preferences can improve LMs via preference optimization, across various model families and scales. Our analyses reveal gaps in cultural awareness among LMs, native speakers, and search engines, especially regarding implicit cultural commonsense. By releasing CARE, we hope to foster the development of more culturally adaptive LMs.
@@
During data collection, we intentionally excluded highly sensitive and controversial topics (e.g., “Are security concerns about the niqab exaggerated or justified?”), as these tend to elicit divergent personal views. Our work addresses generally shared topics within each cultural group (e.g., “Do Chinese people like stinky tofu?” in the Opinion category of CARE), allowing us to capture collective preferences that reflect broader cultural consensus Bakker et al. (2022) with a neutral point of view (e.g., “Some Chinese like stinky tofu. It is a traditional food made by $\ldots$ It is most popular in $\ldots$ regions. At the same time, many Chinese people don’t like it $\ldots$” ). This approach facilitates more objective and replicable evaluation, as evidenced by the high inter-annotator agreement reported in §3.2, thus supports more reliable algorithmic comparisons, which are the primary focus of this paper. The scope of our study is complementary to other work  Chiu et al. (2024a); Kirk et al. (2024); Huang et al. (2025) that explored diverse values and subjective viewpoints Zhong et al. (2021).
{CJK*}
UTF8gbsn We also note that LMs may occasionally respond to culture-specific questions in a stereotypical or biased manner, stemming from a lack of cultural understanding. For instance, when asked “In China, what does the term “laowai” mean when referring to foreigners?”, LMs often incorrectly interpret the term as disrespectful. In reality, “laowai (老外)” is typically a neutral descriptor in Chinese, used to denote foreigners without negative connotations. While CARE includes examples that clarify such misunderstandings through human-written responses, future studies in cultural red-teaming can further investigate and address these failure cases.
Ethics Statement
In this study, we employed in-house annotators to collect cultural samples and provide preference judgments in CARE. The annotators for the Chinese and Arabic samples were university-level students fluent in the respective languages. Japanese annotators were Japanese workers with university degrees and cross-cultural experience abroad. Each Chinese and Arabic annotator was paid at \$18 per hour, exceeding the U.S. federal minimum wage. We ensured that no personal data was collected from the annotators and emphasized that participation was voluntary. We also informed the annotators that the collected data would be used to enhance the cultural awareness of LMs. The culture-related questions and reference responses in CARE are either from the existing works or human-written sentences provided by our annotators.
@@ Acknowledgments
The authors would like to thank Chao Jiang, Jad Matthew Bardawil, Nour Allah El Senary, Ruohao Guo, Xiaofeng Wu, Yuming Pan, and Zirui Shao for their help in data annotation; Yao Dou and Xiaofeng Wu for their helpful feedback. We would also like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This work was funded in part through a Sony Faculty Innovation Award, and by the NSF under grant number IIS-2144493, IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 Annotation Details
To construct the question-answer pairs in CARE, we recruited native Chinese, Arabic, and Japanese speakers as mentioned in § 3.2. Among our annotators, 5 Japanese annotators were workers with experience living abroad (e.g., in the US), while the rest were international college students. Annotators were paid at the rate of 18 USD/hour and told that they were helping create and annotate data for improving the cultural awareness of LMs. The detailed annotation guideline for dataset construction is provided in Figures 15 to 19, and the instruction for preference annotation is provided in Figures 13. We also show our rank-and-rate web interface for collecting human cultural preference pairs in Figure 14.
A.2 Annotation Agreements
To assess annotators’ agreement on this culture preference judgment task, we invite additional native speakers for each culture, and report the Pearson correlation on different cultural categories in Table 11. It shows that high correlation is achieved among annotators across five cultural categories.
Culture Entities Opinion Norms Commonsense Literacy
Arab 0.96 0.90 0.84 0.88 0.97
Chinese 0.84 0.90 0.96 0.94 0.99
Japanese 0.99 0.88 0.96 0.85 0.99
Table 11: Inter-annotator agreement measured by Spearman’s $\rho$ on five cultural categories. Agreement for this preference judgment task is consistently high across different cultural categories.
A.3 Statistics
Table 8 shows the detailed statistics for Chinese, Arab, and Japanese cultures in CARE across the cultural categories, stratified by the data source. The questions specific to Chinese, Arab, and Japanese cultures are written in Chinese, Arabic, and Japanese, respectively. The culturally-relevant samples obtained from the instruction-tuning dataset mostly fall within the cultural entities, cultural opinion, or literacy categories, while samples obtained from cultural knowledge bases provide more social norms and cultural commonsense data.
Table 9 shows the detailed statistics for other foreign cultures in CARE for each cultural category. These samples were obtained when filtering the instruction tuning datasets and cultural knowledge bases (§3.1). We wrote these foreign samples in the native language and used them as part of our training set and in our analysis on the impact of the source culture in preference learning (§4.5). We also manually collected more samples for social norms and cultural commonsense from online websites to cross 100 samples in each category. We ensured these foreign samples do not include anything relevant to Chinese, Arab, or Japanese cultures.
Table 10 shows the detailed statistics for Chinese, Arabic, and Japanese samples in CARE across the geographic scope. Most samples fall into nationwide category, while sub-nationwide samples provide insights about more detailed cultural information, continent-wide and worldwide samples assess understanding between several cultures.
A.4 Comparison with existing datasets
Table [Uncaptioned image]CARE: Assessing the Impact of Multilingual Human Preference Learning on Cultural Awareness presents illustrative examples from representative datasets that aim to improve LMs’ cultural awareness. While CultureLLM Li et al. (2024a) focuses on synthetic data without preference supervision, and Prism Kirk et al. (2024) provides human ratings on generic English questions, neither provides proper multilingual and culture-specific contexts with human preferences. In contrast, CARE provides human preferences on culturally grounded topics across multiple languages.
@@ Appendix B Implementation Details
SFT Tuning.
We conduct full-parameter SFT on the instruction data only, using the questions and human-written reference responses in CARE. We tune the learning rate in the range $\{1e^{-5},2e^{-5}\}$, and train with a batch size of 128 on 4–8 NVIDIA A40 GPUs.
Preference Tuning.
We perform full-parameter preference optimization individually for each language using DPO, KTO, and SimPO for 3 epochs until the loss converges. Training is done with a batch size of 128 on 4-8 NVIDIA A40 GPUs. We tune the learning rate in the range $\{3e^{-7},5e^{-7},7e^{-7}\}$ and set beta as 0.1. The training involves full fine-tuning with 5 warmup steps and employs a linear learning rate scheduler.
MAPO Tuning.
We use the same base LLM to be tuned as the rollout model and use NLLB-600M-distilled Costa-Jussà et al. (2022) to calculate the alignment score as suggested in the original paper She et al. (2024). We conduct DPO tuning on the generated preference pairs and report the results of the first iteration.
@@
For open-sourced LMs, we run inference on one NVIDIA A40 GPU with the vLLM library [^2] https://docs.vllm.ai Kwon et al. (2023). We perform decoding by setting the following parameters {temperature=0.7, top_p=1}. We limit the context length by setting {max_model_len=2048}. We also limit the number of generated tokens by the models by setting {max_tokens=1024}. For the closed-source GPT-4o LM, we run inference with Azure OpenAI API.
Baseline preference datasets.
Table 12 lists the URLs of the multilingual adaptations of the general-domain preference datasets used in our experiments (§4.2). HelpSteer3 is natively multilingual, with prompts and responses originally written in various languages and manually annotated preference labels. In contrast, the multilingual versions of UltraFeedback and OpenOrca were generated by automatically translating the English response pairs into the target languages, followed by automatic quality control to remove poorly translated samples. Their preference signals are also synthetic. UltraFeedback relies on GPT-4 to rank the responses, while OpenOrca assumes GPT-4’s responses are always superior to Llama’s when constructing preference pairs  Wang et al. (2025a).
Dataset Language Source HuggingFace Repository
OpenOrca ar translation 2A2I/argilla-dpo-mix-7k-arabic
ja translation yongtae-jp/orca_dpo_pairs_ja
zh translation wenbopan/Chinese-dpo-pairs
UltraFeedback ar translation 2A2I/argilla-dpo-mix-7k-arabic
zh translation wenbopan/Chinese-dpo-pairs
HelpSteer3 zh, ja native nvidia/HelpSteer3
Table 12: Multilingual adaptations of baseline preference datasets. “Source” indicates whether the non-English data are natively collected or derived via automatic translation.
@@
You are a native [Chinese/Arab/Japanese] person, familiar
with [Chinese/Arab/Japanese] culture and traditions.
@@
Figure 7: Prompt template for role-play inference. The LM is told to take on the persona of a native [Chinese/Arab/Japanese] person who is familiar with the culture’s traditions, and then asked the culture-specific {Question}.
@@
Chinese Arabic Japanese Model Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide [Uncaptioned image] Gemma2-27B 6.50 7.16 8.19 6.47 4.71 6.73 4.88 6.29 2.50 5.89 7.20 [Uncaptioned image] Llama3.3-70B 6.20 6.92 7.68 5.57 5.14 6.16 4.31 5.50 6.75 5.77 6.40 [Uncaptioned image] Qwen2.5-72B 7.70 8.49 8.60 6.78 7.86 7.62 5.19 8.58 4.25 6.57 8.00 [Uncaptioned image] Mistral-Large 7.42 8.11 8.73 6.70 5.29 7.24 5.31 8.08 4.50 6.45 6.40 [Uncaptioned image] GPT-4o 8.59 8.74 8.72 7.11 7.43 8.15 7.79 8.92 7.00 8.32 8.00
Table 13: Performance comparison w.r.t. geographic scope. Scores are computed on the entire CARE data, including both local culture and foreign culture.
@@ Appendix D Additional Results
D.1 Performance by Preference Learning strategies
Table C.2 shows per-category performance when tuned with different preference learning strategies.
D.2 Performance by Geographic Scope
Table 13 shows results from different LMs when stratified by geographic scope. We find that the models struggle the most with sub-nationwide questions. Most models achieve the best scores on continent-wide or worldwide questions.
D.3 Performance on general NLP tasks
We then examine whether cultural preference learning impacts the model’s overall knowledge and capabilities, using well-established benchmarks for Chinese, Arabic, Japanese, and English: ArabicMMLU Koto et al. (2024), ChineseMMLU Li et al. (2023), the Japanese subset of MMLU-ProX Xuan et al. (2025), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), and WinoGender Rudinger et al. (2018). We compare different versions of Llama3.1-8B-Instruct: the vanilla LM, the aligned LM with general human preference from the combined UltraFeedback and OpenOrca datasets, and the culturally aligned LM with cultural human preference from CARE.
The results in Table 15 show very small differences between the vanilla LM and its aligned version, on benchmarks in both native language and English, suggesting that cultural preference learning does not hinder the model’s overall capabilities.
Arabic Chinese Japanese Model Vanilla DPO (w. Orca+Ultra) DPO (w. CARE) Vanilla DPO (w. Orca+Ultra) DPO (w. CARE) Vanilla DPO (w. Orca) DPO (w. CARE) MMLU (ar/zh/ja) 56.4 57.1 56.5 55.5 55.5 55.2 28.8 28.4 30.2 MMLU (en) 69.1 68.8 69.1 69.1 68.9 69.1 69.1 68.8 69.0 TruthfulQA 37.8 38.7 38.6 37.8 38.8 40.8 37.8 41.1 38.7 WinoGender 52.6 55.0 52.8 52.6 54.4 52.1 52.6 54.7 52.9
Table 15: Accuracy of Llama3.1-8B-Instruct variants on general benchmarks assessing overall capabilities in native languages (Chinese, Arabic, and Japanese) and English. Slight differences are observed among the LM variants after applying different preference learning methods.
@@ Appendix E Responsible NLP Checklist
Our annotators were undergraduate student employees paid at their normal hourly rate of 18 USD per hour. All annotators were informed that they were assessing data to be used in a research project focused on “studying the cultural awareness of various language models”.